# Measuring cosmic baryon density with FRB and GW data

### v1

@@ 4 Conclusion
In this Letter, we combine two emerging cosmological probes, i.e., $104$ localized FRBs and $47$ GW standard sirens, to measure the cosmic baryon density in a robust manner. DM-based FRB cosmology struggles with various parameter degeneracies, particularly between $H_{0}$ and $\Omega_{\rm b}$. GWs can independently infer $H_{0}$ from self-calibrated luminosity distances obtained through waveform analysis as sirens, thus breaking the $\Omega_{\rm b}$-$H_{0}$ degeneracy. This joint analysis yields a constraint of $\Omega_{\rm b}=0.0488\pm 0.0064$ ($1\sigma$), which is remarkably consistent with early-Universe values from BBN and CMB at the $\sim 13\%$ level and provides a purely late-Universe perspective on resolving the missing baryon problem. We also obtain a simultaneous constraint on the Hubble constant, $H_{0}=71.6^{+4.4}_{-8.6}~{}\mathrm{km\,s^{-1}\,Mpc^{-1}}$, primarily driven by the GW data. Together, these results exhibit $\Omega_{\rm b}$-$H_{0}$ concordance between the early and late Universe.
@@
Moreover, we explore two prior-based scenarios: (i) $H_{0}$-prior-based cases using FRB data alone, where adopting Gaussian priors from Planck18 and SH0ES yields $\Omega_{\rm b}=0.0513\pm 0.0046$ and $\Omega_{\rm b}=0.0473\pm 0.0043$, respectively. These results indicate that the $H_{0}$ tension would bias baryon density estimates, although the induced “$\Omega_{\rm b}~{}$inconsistency” remains within $1\sigma$ uncertainty in our sample. (ii) Simulation-based prior cases using FRB + GW, where fixing empirical DM model parameters from the IllustrisTNG simulation and further fixing $f_{\rm d}$ to a redshift-dependent form yield $\Omega_{\rm b}=0.0491\pm 0.0054$ and $0.0507\pm 0.0052$, respectively. These two results agree with that of FRB + GW at the $1\sigma$ level, indicating that simulation-based priors can improve current constraints without introducing significant bias. However, in both (i) and (ii), systematic effects from external assumptions (related to $H_{0}$, $f_{\rm d}$, and empirical DM model parameters) may become non-negligible when applied to larger future FRB and GW samples.
@@
Overall, the novel combination of FRBs and GWs enables a direct, unique, and robust measurement of both the baryon content and cosmological distances at low redshifts. With an increasing number of GW detections expected in the upcoming LVK run, and potentially thousands of localized FRBs expected from CHIME/Outriggers (Masui, 2023) and DSA-110 (Law et al., 2024), their synergy will soon emerge as a powerful low-redshift cosmological probe.
5 Acknowledgments
We thank Tian-Nuo Li, Meng-Lin Zhao, Yang Liu, and Qin Wu for fruitful discussions. This work was supported by the National SKA Program of China (grant Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (grant Nos. 12473001, 11975072, 11875102, and 11835009), the China Manned Space Program (grant No. CMS-CSST-2025-A02), and the National 111 Project (grant No. B16009).
@@ Appendix A Constraints on $H_{0}$ and DM parameters
Figure 3: Same as Figure 1 but for posterior distribution of $H_{0}$. The median and $1\sigma$ credible interval from FRB + GW is $H_{0}=71.6^{+4.4}_{-8.6}$ km s-1  Mpc-1. The orange and black errorbars correspond to constraints from FRB + $\Omega_{\rm b}h^{2}$ (BBN prior (Cooke et al., 2018)) and GW data alone, respectively. The green, yellow, red, and gray shaded bands denote the results from Planck18 (Aghanim et al., 2020), SH0ES (Riess et al., 2022)), Wu22 (Wu et al., 2022), and James22 (James et al., 2022), respectively. Note that Wu22 adopts the TNG prior, and both Wu22 and James22 adopt $\Omega_{\rm b}h^{2}$ priors from BBN (Cooke et al., 2018) and CMB (Aghanim et al., 2020), respectively. All results are at $1\sigma$ uncertainty.
@@
The constraints on $H_{0}$ are listed in Table 1, and posteriors and corresponding credible intervals for $H_{0}$ are shown in Figure 3. A direct measurement of $H_{0}$ using FRBs alone is difficult due to strong parameter degeneracies. Specifically, our measurement using FRBs alone is $H_{0}=77.0^{+20.0}_{-40.0}$ km s-1  Mpc-1, and the inclusion of GW data narrows the constraint to $H_{0}=71.6^{+4.4}_{-8.6}$ km s-1  Mpc-1. This improvement is mainly driven by the GW data itself, which yields $H_{0}=72.6^{+5.2}_{-8.9}$ km s-1  Mpc-1. Previous studies predominantly inferred $H_{0}$ by fixing the baryon density parameter at the early-Universe value. For example, Wu et al. (2022) (hereafter Wu22) and James et al. (2022) (hereafter James22) used different FRB samples to derive $H_{0}=69.31^{+6.21}_{-6.63}~{}\mathrm{km~{}s^{-1}~{}Mpc^{-1}}$ and $H_{0}=73^{+12}_{-8}~{}\mathrm{km~{}s^{-1}~{}Mpc^{-1}}$, by assuming external priors on $\Omega_{\rm b}h^{2}\sim\mathcal{N}(0.02235,0.00049)$ from BBN and $\Omega_{\rm b}h^{2}=0.02242$ from Planck18, respectively, which are shown as red and gray bands in Figure 3. Recently, Wang et al. (2025) and Gao et al. (2025a) also used $92$ and $108$ localized FRBs to obtain $H_{0}=69.04_{-2.07}^{+2.30}\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$ and $H_{0}=69.40_{-1.97}^{+2.14}\mathrm{~{}km}\mathrm{~{}s}^{-1}\mathrm{Mpc}^{-1}$, respectively, both assuming BBN and TNG priors. Using the same BBN prior, we obtain $H_{0}=66.2^{+5.3}_{-7.3}$ km s-1  Mpc-1from FRB + $\Omega_{\rm b}h^{2}$, which slightly favors the Planck18 result with a precision comparable to that of FRB + GW. Unlike prior-based methods, the FRB + GW constraint provides a late-Universe result of $H_{0}$. More data is required to offer further insight into arbitrating the $H_{0}$ tension. Notably, time-delay cosmography using strongly lensed FRBs by massive galaxies can offer an independent measurement of $H_{0}$, since the time delays between lensed FRB images can be determined with exceptional precision (Li et al., 2018; Zhang et al., 2025a).
For DM parameters, the constraints are $f_{\rm d}=0.90\pm 0.04$, $F=0.36^{+0.13}_{-0.05}$, $\mu_{\rm host}=4.67^{+0.20}_{-0.16}$, and $\sigma_{\rm host}=0.95^{+0.09}_{-0.13}$. The inclusion of GW data has a negligible impact on this constraint. For the host contribution, the results of $\mu_{\rm host}$ and $\sigma_{\rm host}$ correspond to a median rest-frame host DM of $108\pm 20~{}\mathrm{pc~{}cm^{-3}}$ (corresponding mean DM is $168^{+57}_{-40}~{}\mathrm{pc~{}cm^{-3}}$), which is consistent with both $\mathrm{DM}_{\rm host}=103_{-48}^{+68}~{}\mathrm{pc~{}cm^{-3}}$ (Yang et al., 2025) and $\mathrm{DM}_{\rm host}=130_{-23}^{+25}~{}\mathrm{pc~{}cm^{-3}}$ (Connor et al., 2025).
@@ Appendix B FRB data
In this Appendix, we list the properties of the $104$ localized FRBs used in this work in Table 2.
[…]
Table 2: Properties of $104$ localized FRBs.
[…]

### v2

@@ IV Conclusion
In this paper, we combine two emerging cosmological probes, i.e., $104$ localized FRBs and $47$ GW standard sirens, within a unified framework to measure the cosmic baryon density in a $H_{0}$-free approach. Most previous baryon censuses assumed a fixed value of $H_{0}$ or adopted an external prior, which could bias the inference in the presence of the Hubble tension. If FRB DMs are instead used to trace the baryons, additional parameter degeneracies arise in the DM modeling, particularly between $H_{0}$ and $\Omega_{\rm b}$. GWs independently infer $H_{0}$ from self-calibrated luminosity distances obtained through waveform analysis as sirens, thus breaking the $\Omega_{\rm b}$-$H_{0}$ degeneracy. This joint analysis yields a constraint of $\Omega_{\rm b}=0.0488\pm 0.0064$ ($1\sigma$), which is remarkably consistent with early-Universe values from BBN and CMB at the $\sim 13\%$ level and provides a purely late-Universe perspective on resolving the missing baryon problem. We also obtain a simultaneous constraint on the Hubble constant, $H_{0}=71.6^{+4.4}_{-8.6}~\mathrm{km\,s^{-1}\,Mpc^{-1}}$, primarily driven by the GW data. Together, these results exhibit $\Omega_{\rm b}$-$H_{0}$ concordance between the early and late Universe.
@@
Moreover, we explore two prior-based scenarios: (i) $H_{0}$-prior-based cases using FRB data alone, where adopting Gaussian priors from Planck18 and SH0ES yields $\Omega_{\rm b}=0.0513\pm 0.0046$ and $\Omega_{\rm b}=0.0473\pm 0.0043$, respectively. These results indicate that the $H_{0}$ tension would bias baryon density estimates, although the induced $\Omega_{\rm b}$~inconsistency” remains within $1\sigma$ uncertainty in our sample. (ii) Simulation-based prior cases using FRB + GW, where fixing empirical DM model parameters from the IllustrisTNG simulation and further fixing $f_{\rm d}$ to a redshift-dependent form yield $\Omega_{\rm b}=0.0491\pm 0.0054$ and $0.0507\pm 0.0052$, respectively. These two results agree with that of FRB + GW at the $1\sigma$ level, indicating that simulation-based priors can improve current constraints without introducing significant bias. However, in both (i) and (ii), systematic effects from external assumptions (related to $H_{0}$, $f_{\rm d}$, and empirical DM model parameters) may become non-negligible when applied to larger future FRB and GW samples.
@@
We caveat that our baryon census still depends on the adopted $\Omega_{\rm m}$ prior and the assumed GW population model. Nevertheless, the $\Omega_{\rm b}$ values inferred when removing these assumptions show only modest shifts and remain consistent with CMB + BBN values at the $1\sigma$ level. Using larger samples of localized FRBs and GW standard sirens, future analyses would jointly constrain $\Omega_{\rm m}$, GW population parameters, and $\Omega_{\rm b}$ more robustly, thereby reducing this dependence.
Overall, the novel combination of FRBs and GWs enables a direct, unique, and robust measurement of both the baryon content and cosmological distances at low redshifts. With an increasing number of GW detections expected in the upcoming LVK run, and potentially thousands of localized FRBs expected from CHIME/Outriggers masui2023pinpointing and DSA-110 Law:2023ibd, their synergy will soon emerge as a powerful low-redshift cosmological probe.
V Acknowledgments
We thank Tian-Nuo Li, Meng-Lin Zhao, Yang Liu, and Qin Wu for fruitful discussions. This work was supported by the National Natural Science Foundation of China (Grants Nos. 12533001, 12575049, 12473001, and 12473091), the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the China Manned Space Program (Grant No. CMS-CSST-2025-A02), the National 111 Project (Grant No. B16009), and the Fundamental Research Funds for the Central Universities (Grant No. N2405008).
@@ Appendix A Constraints on $H_{0}$ and DM parameters
Figure 4: Same as Fig. 1 but for posterior distribution of $H_{0}$. The median and $1\sigma$ credible interval from FRB + GW is $H_{0}=71.6^{+4.4}_{-8.6}$ km s-1  Mpc-1. The orange and black errorbars correspond to constraints from FRB + $\Omega_{\rm b}h^{2}$ (BBN prior Cooke:2017cwo) and GW data alone, respectively. The green, yellow, red, and gray shaded bands denote the results from Planck18 2020Planck, SH0ES Riess:2021jrx), Wu22 Wu:2021jyk, and James22 James:2022dcx, respectively. Note that Wu22 adopts the TNG prior, and both Wu22 and James22 adopt $\Omega_{\rm b}h^{2}$ priors from BBN Cooke:2017cwo and CMB 2020Planck, respectively. All results are at $1\sigma$ uncertainty.
@@
The constraints on $H_{0}$ are listed in Table 1, and posteriors and corresponding credible intervals for $H_{0}$ are shown in Fig. 4. A direct measurement of $H_{0}$ using FRBs alone is difficult due to strong parameter degeneracies. Specifically, our measurement using FRBs alone is $H_{0}=77.0^{+20.0}_{-40.0}$ km s-1  Mpc-1, and the inclusion of GW data narrows the constraint to $H_{0}=71.6^{+4.4}_{-8.6}$ km s-1  Mpc-1. This improvement is mainly driven by the GW data itself, which yields $H_{0}=72.6^{+5.2}_{-8.9}$ km s-1  Mpc-1. Previous studies predominantly inferred $H_{0}$ by fixing the baryon density parameter at the early-Universe value. For example, Refs. Wu:2021jyk (hereafter Wu22) and James:2022dcx (hereafter James22) used different FRB samples to derive $H_{0}=69.31^{+6.21}_{-6.63}~\mathrm{km~s^{-1}~Mpc^{-1}}$ and $H_{0}=73^{+12}_{-8}~\mathrm{km~s^{-1}~Mpc^{-1}}$, by assuming external priors on $\Omega_{\rm b}h^{2}\sim\mathcal{N}(0.02235,0.00049)$ from BBN and $\Omega_{\rm b}h^{2}=0.02242$ from Planck18, respectively, which are shown as red and gray bands in Fig. 4. Recently, Refs. Wang:2025ugc and Gao:2025fcr also used $92$ and $108$ localized FRBs to obtain $H_{0}=69.04_{-2.07}^{+2.30}\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$ and $H_{0}=69.40_{-1.97}^{+2.14}\mathrm{~km}\mathrm{~s}^{-1}\mathrm{Mpc}^{-1}$, respectively, both assuming BBN and TNG priors. Using the same BBN prior, we obtain $H_{0}=66.2^{+5.3}_{-7.3}$ km s-1  Mpc-1from FRB + $\Omega_{\rm b}h^{2}$, which slightly favors the Planck18 result with a precision comparable to that of FRB + GW. Unlike prior-based methods, the FRB + GW constraint provides a late-Universe result of $H_{0}$. More data is required to offer further insight into arbitrating the $H_{0}$ tension. Notably, time-delay cosmography using strongly lensed FRBs by massive galaxies can offer an independent measurement of $H_{0}$, since the time delays between lensed FRB images can be determined with exceptional precision Li:2017mek; Zhang:2024rra.
For DM parameters, the constraints are $f_{\rm d}=0.90\pm 0.04$, $F=0.36^{+0.13}_{-0.05}$, $\mu_{\rm host}=4.67^{+0.20}_{-0.16}$, and $\sigma_{\rm host}=0.95^{+0.09}_{-0.13}$. The inclusion of GW data has a negligible impact on this constraint. For the host contribution, the results of $\mu_{\rm host}$ and $\sigma_{\rm host}$ correspond to a median rest-frame host DM of $108\pm 20~\mathrm{pc~cm^{-3}}$ (corresponding mean DM is $168^{+57}_{-40}~\mathrm{pc~cm^{-3}}$), which is consistent with both $\mathrm{DM}_{\rm host}=103_{-48}^{+68}~\mathrm{pc~cm^{-3}}$ Yang:2024vqq and $\mathrm{DM}_{\rm host}=130_{-23}^{+25}~\mathrm{pc~cm^{-3}}$ connor2025gas.
@@ Appendix B FRB data
In this Appendix, we list the properties of the $104$ localized FRBs used in this work in Table LABEL:tab:localized_FRBs.
Table 2: Properties of $104$ localized FRBs. The listed parameters include redshift, right ascension (R.A.), declination (Decl.), observed dispersion measure (DM), the Milky Way ISM contribution derived from the NE2001 model, and types of their host galaxies. The host types are classified into three categories: type 1 refers to repeating FRBs in dwarf galaxies, type 2 to repeating FRBs in spiral galaxies, and type 3 to one-off FRBs.
Name Redshift $\mathrm{DM}_{\mathrm{obs}}$ $\mathrm{DM}_{\mathrm{MW,ISM}}$ (NE2001) R.A. (J2000) Decl. (J2000) Host Type Reference